OpenStudy (anonymous):

Convert 3cis 180° to rectangular form.

OpenStudy (anonymous):

do you know \(\cos(180^\circ)\) ?

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

ok how about \(\sin(180^\circ)\)?

OpenStudy (anonymous):

o

OpenStudy (anonymous):

k then you are pretty much done \[3(\cos(180)+i\sin(180))=?\]

OpenStudy (anonymous):

-3i

OpenStudy (anonymous):

i think it is -3